What is the energy levels for zAz-1
1.En forAz-1=Z2*En for H
2.En for AZ-1=Z*En for H

Grade:

1 Answers

Arun
5 years ago
energy levels for zA(+z1)− can be given by (1) En for zA(+z–1) = Z2 × En for H (2) En for zA(+z–1) = Z× En for H (3) En for zA(+z–1) = 12Z × En for H (4) En for zA(+z–1)= 1 Z × En for H The specific charge for positive rays is much less than the specific charge for cathode rays. This is because (1) Possible rays are positively ...